### v1

# FakeFormer: Efficient Vulnerability-Driven Transformers for Generalisable Deepfake Detection

Abstract
Recently, Vision Transformers (ViTs) have achieved unprecedented effectiveness in the general domain of image classification. Nonetheless, these models remain underexplored in the field of deepfake detection, given their lower performance as compared to Convolution Neural Networks (CNNs) in that specific context. In this paper, we start by investigating why plain ViT architectures exhibit a suboptimal performance when dealing with the detection of facial forgeries. Our analysis reveals that, as compared to CNNs, ViT struggles to model localized forgery artifacts that typically characterize deepfakes. Based on this observation, we propose a deepfake detection framework called FakeFormer, which extends ViTs to enforce the extraction of subtle inconsistency-prone information. For that purpose, an explicit attention learning guided by artifact-vulnerable patches and tailored to ViTs is introduced. Extensive experiments are conducted on diverse well-known datasets, including FF++, Celeb-DF, WildDeepfake, DFD, DFDCP, and DFDC. The results show that FakeFormer outperforms the state-of-the-art in terms of generalization and computational cost, without the need for large-scale training datasets. The code is available at https://github.com/10Ring/FakeFormer.
1 Introduction
Figure 1: Comparison of FakeFormer (and FakeSwin) to existing methods including SBI [47], CADDM [21], and Transformer-based approaches [60], namely TALL-Swin and ViT-B+TALL, in terms of model size, FLOPs, and AUC. The size of each bubble represents the number of model parameters. All methods are trained on FF++ [46] and tested on CDF2 [38].
In the era of deep learning, it has never been easier to generate ultra-realistic facial images, undetectable to the naked eye. Such forged visual data, commonly called deepfakes [48, 31], are produced by manipulating the identity or the expression of a subject using generative deep neural networks such as Generative Adversarial Networks (GAN) [25]. Although beneficial in several applications, deepfakes can also be harmful to society if employed for malicious purposes such as spreading misinformation and fraud.
To face this growing threat, the research community is making ceaseless efforts to introduce approaches for automatically detecting deepfakes. Currently, the most successful deepfake detection methods are mostly relying on Convolution Neural Networks (CNNs) [21, 28, 36, 58, 47, 63, 11, 45, 13, 12, 62, 4, 3, 2]. Paradoxically, thanks to their simplicity and scalability, Vision Transformers [23, 53] (ViTs) are becoming increasingly popular in other visual classification tasks, often outperforming CNN architectures. However, ViT architectures remain relatively underexplored in the field of deepfake detection, given their lower performance as compared to CNNs in that context [22, 55].
In this paper, we aim to: (1) investigate the reasons behind this drop in performance with respect to CNNs and (2) suggest accordingly adequate yet simple strategies for enhancing the performance of ViT architectures in the context of deepfake detection, while benefiting from their efficiency. Specifically, we first experimentally analyze the capability of ViTs to capture localized artifacts typically characterizing deepfakes, as compared to CNNs. Our analysis suggests that ViTs struggle more in this regard because they tend to focus on global context [23, 15, 29, 40, 56].
Therefore, to address this issue, we propose to inject a local attention mechanism in ViTs by explicitly detecting vulnerable patches using blending-based data synthesis techniques such as [47, 37]. For that purpose, we extend the notion of vulnerable points [45], which represent the pixels that are the most impacted by the blending operation, to vulnerable patches. These patches are the group of pixels that are the most likely to incorporate blending artifacts and can be used in conjunction with any transformer-like architecture. This attention module is referred to as Learning-based Local Attention mechanism (L2-Att). Even though the idea of attention to vulnerabilities has been explored in [45], our experimental study shows that generalizing this approach to transformers contributes to:
  1. 1.
    Enhanced generalization performance: As reflected in Figure 1, higher generalization capabilities are achieved by different transformer-based architectures enclosing the proposed L2-Att strategy, namely FakeFormer and FakeSwin. Specifically, the FakeSwin is a variant of FakeFormer where Swin [40] is used as a backbone. The notations -S and -B refer to the backbone size (small and base, respectively)111 More details are reported in Section 5..
  2. 2.
    Friendly computational cost and reduced model size: As illustrated in Figure 1, although achieving impressive performance, the proposed transformer models cost a relatively low number of FLOPs, and necessitate fewer parameters as compared to other methods.
  3. 3.
    Mitigating the need for large-scale datasets: ViTs are known to be data-hungry [23, 61, 10, 53]. Therefore, by employing generic blending-based synthesis methods as [37, 47] to define pseudo-vulnerable patches, we mitigate this problem.
This is confirmed by extensive experiments conducted on seven challenging datasets including FF++ [46], two versions of CDF [38], DFD [24], DFW [65], DFDCP [20], and DFDC [19]. Our method outperforms state-of-the-art approaches, including both CNN-based and ViT-based methods [22, 33, 55], although the latter [22, 33, 55] usually rely on a significantly larger training set.
Paper Organization. Section 2 discusses related work. Section 3 analyzes the performance gap between ViTs and CNNs in deepfake detection. The proposed approach called FakeFormer is presented in Section 4. Section 5 and Section 6 discuss the experiments and the limitations, respectively. Section 7 concludes this work and suggests future directions.
Figure 2: Examples are randomly selected to illustrate the four types of deepfakes in common FF++ [46] dataset. It can be observed that Face2Face [52] and NeuralTextures [51] exhibit more subtle artifacts.
[…]
2 Related Work
Deepfake Detection with CNNs.
In the field of deepfake detection, Convolution Neural Networks (CNNs) are often favored over other types of architectures. Pioneer works in deepfake detection have predominantly employed established CNN architectures such as XceptionNet [14] as binary classifiers [46, 1, 6]. However, these methods struggle to generalize to unseen manipulation methods. To address this challenge, a wide range of strategies have been investigated [21, 36, 7, 62, 47, 37, 43] such as disentanglement learning [7, 62], multi-task learning [11, 45, 63], and pseudo-fake synthesis [47, 37, 63, 12, 43]. With the advances in generative modeling, deepfakes are constantly gaining in realism, resulting in invisible localized artifacts. The aforementioned methods are exposed to the risk of becoming obsolete. Indeed, they still rely on standard CNNs, consequently facing a loss of local information through successive convolution layers [64, 57]. For capturing more effectively low-level features, methods with implicit attention strategies have been proposed [64, 57, 54], but have shown poor generalization [45]. Recently, Nguyen et al. [45] argued that to ensure robustness to high-quality deepfakes while maintaining good generalization capabilities, an explicit attention mechanism within a multi-task learning framework is needed. In particular, they introduce LAA-Net, a network guided to focus on vulnerable points (defined as the pixels that are the most likely to incorporate blending artifacts). To simulate vulnerable pixels, blending-based data synthesis techniques are leveraged. However, although LAA-Net achieves state-of-the-art performances, it is sensitive to noise and is based on a cumbersome CNN that is computationally costly, thereby limiting its practical deployment. In this paper, we investigate how to adapt such an explicit attention mechanism to ViT-based architectures to improve the effectiveness of transformer-based deepfake detection methods and benefit from more efficient models.
Deepfake Detection with ViTs. Vision Transformers (ViTs) have attracted the great interest of the research community, as they are claimed to outperform CNNs in terms of both effectiveness and efficiency. While such a statement holds for several applications [50, 23, 61, 30, 53, 8], it does not apply to the field of deepfake detection. So far, only a handful of transformer-based deepfake detection methods have been explored in the literature. For instance, ICT [22] utilizes a basic ViT to detect deepfakes by assessing identity consistency between inner and outer faces in suspect images. To capture multi-scale features, M2TR [55] and DFDT [33] simultaneously process inputs through multiple transformers with different patch sizes. However, despite training on large-scale datasets [27] and/or requiring costly computational resources, these methods still exhibit limited generalization performance compared to CNN-based approaches. This motivates us to investigate in Section 3 the underlying reasons explaining the unsuitability of ViTs for this use-case.
Figure 4: The proposed method: (a) the overall FakeFormer framework, (b) the L2-Att module, and (c) the generation of vulnerable patches.
3 Vision Transformer and Deepfake Detection: Where is the Gap?
In this section, we investigate the specific challenges associated with the use of plain ViTs in deepfake detection. Our primary hypothesis is as follows: Compared to CNNs, ViTs focus on global image representations [23, 15, 29, 40, 56], given their patch-based architecture. Consequently, they struggle to effectively capture local features [15, 29] that are crucial for identifying subtle artifacts in deepfakes. As a result, CNN-based architectures often outperform ViT-based ones in this context. The plausibility of this assumption is investigated by conducting the experiments described below.
Performance with respect to the quality of deepfakes. Here, our goal is to quantify the performance of ViTs and CNNs with respect to the quality of the encountered deepfakes. To that aim, we compare in Figure 3(a) the performance of CNN-based methods (LAA-Net [45], CADDM [21], and SBI [47]) to transformer-based approaches (Plain ViT [23] and Swin [40]) [^2] Note that the small(-S) architecture configuration is employed for both ViT and Swin. The FLOPs and the number of parameters are reported in Section 5. across different Mask Self-Similarity Index Measure (SSIM) [42] ranges on the Celeb-DF [38] dataset. The Mask SSIM has been commonly used for assessing the quality of deepfakes [38, 55, 45] and is obtained by calculating the SSIM [59] score in the facial region between the fake and the corresponding original images. Higher Mask SSIM values indicate a higher perceptual quality of deepfakes [38, 55, 5, 45], generally implying the presence of more localized artifacts. It can be noted from Figure 3(a) that the performance of ViT drops more significantly for higher Mask SSIM values as compared to most CNN-based methods (LAA-Net and CADDM). The performance of Swin, on the other hand, does not deteriorate as much due to its ability to extract low-level features with its local window design; however, it still shows less stability than LAA-Net. These observations support our hypothesis. We specify that for a fair comparison with the state-of-the-art (SOTA), we train ViT and Swin using the same data synthesis [47] used in SBI and LAA-Net.
[…]
As F2F and NT correspond to face reenactment manipulations while FS and DF represent face-swap approaches, it is more likely to observe more subtle artifacts in F2F and NTs. When comparing ViT112p16 and ViT112p8, it can be seen that a ViT with smaller patches exhibits faster convergence compared to those with larger ones. Moreover, increasing the input resolution while conserving the same patch size (see ViT112p8 and ViT224p8) contributes to the amplification of the local information encoded in each patch, also resulting in faster convergence. In both cases, this convergence gap is even more pronounced for more subtle deepfake types such as F2F and NT, indicating the importance of locality in detecting deepfakes with subtle inconsistencies. However, reducing the patch size leads to a quadratically increasing complexity (i.e. the FLOPs are $2.2$G, $8.4$G, and $33.6$G for ViT112p16, ViT112p8, and ViT224p8, respectively). Additionally, it is to note that CNNs converge more rapidly compared to ViTs under different setups (i.e., even with the smallest patch size). This also highlights the fact that CNNs can extract local features more effectively.
Hence, we posit that by proposing a mechanism that allows focusing on subtle artifact-prone regions, we can enhance the performance of ViTs for the task of deepfake detection. While some attempts have been made to introduce local ViTs such as Swin [40], we argue that this remains insufficient for effectively detecting deepfakes. As demonstrated for CNNs [45], incorporating local features does not guarantee that artifact-prone are considered, highlighting the need to introduce attention strategies for explicitly focusing on localized artifacts.
4 FakeFormer
Based on the observations made in Section 3, we propose FakeFormer, a Transformer-based approach integrating an attention mechanism that induces the modeling of subtle inconsistencies. Aside from its effectiveness, FakeFormer is not data-hungry and is computationally efficient. An overview of the proposed approach can be seen in Figure 4, consisting of a Vision Transformer (ViT) coupled with a Learning-based Local Attention module (L2-Att). L2-Att predicts the locations of vulnerable patches which generalize the notion of vulnerable points introduced in [45]. FakeFormer is trained using only normal data and utilizes blending-based data synthesis techniques [37, 47]. It can be noted that we utilize ViT to formulate our methodology for the sake of simplicity. However, our method is compatible with other transformer-based architectures, as demonstrated in Section 5. In what follows, we depict the different components of FakeFormer.
4.1 Vision Transformer (ViT)
Given an image $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ as input, we first reshape it into a sequence of non-overlapping flattened 2D patches, denoted as $\{\mathbf{x}_{i}\in\mathbb{R}^{C.P^{2}}\text{ with }i\in[\![1,N]\!]\}$, where $(H,W)$ represents the input resolution, $C$ denotes the number of channels, $P\times P$ indicates the size of an image patch, and $N=\frac{H\times W}{P^{2}}$ denotes the number of patches. The ViT linearly maps each $\mathbf{x}_{i}$ into a patch embedding $\mathbf{z}^{0}_{i}\in\mathbb{R}^{D}$ using a learnable matrix $\mathbf{E}\in\mathbb{R}^{(C.P^{2})\times D}$. Subsequently, a learnable embedding $\mathbf{x}_{cls}\in\mathbb{R}^{D}$ is prepended at the zero-index of embeddings $\mathbf{z}^{0}$ for the classification. Additionally, we use a learnable positional embedding $\mathbf{E}^{pos}$ to incorporate the position information of patches. The aforementioned process is described as follows,
$\mathbf{z}^{0}=[\mathbf{x}_{cls};\mathbf{x}_{1}\mathbf{E};\mathbf{x}_{2}\mathbf{E};\cdots;\mathbf{x}_{N}\mathbf{E}]+\mathbf{E}^{pos},$ (1)
where $\mathbf{E}^{pos}\in\mathbb{R}^{(N+1)\times D}$ and $\mathbf{z}^{0}\in\mathbb{R}^{(N+1)\times D}$. Afterward, $\mathbf{z}^{0}$ is fed into several transformer encoder blocks. Similar to ViT [23], FakeFormer has $L$ blocks, each one containing a multi-head self-attention (MHSA), Layernorm (LN), and a multi-layer perceptron (MLP). The feature extraction process is described as follows,
$\mathbf{z}^{l}$ $=\text{MHSA}(\text{LN}(\mathbf{z}^{l-1}))+\mathbf{z}^{l-1},$ (2)
$\mathbf{z}^{l^{\prime}}$ $=\text{MLP}(\text{LN}(\mathbf{z}^{l}))+\mathbf{z}^{l},$
with $l\in[\![1,L]\!]$ and $\mathbf{z}^{l},\mathbf{z}^{l^{\prime}}\in\mathbb{R}^{(N+1)\times D}$. The extracted feature from the classification embedding $\mathbf{z}^{L^{\prime}}_{0}$ is processed by a classification head composed of an MLP, resulting in the predicted category output $\hat{\mathbf{y}}$. In the task of deepfake detection, the categories consist of real or fake. Gray blocks in Figure 4a are parts of the used ViT.
4.2 Attention to Vulnerable Patches
To overcome the challenge discussed in Section 3, we propose integrating the Learning-based Local Attention Module (L2-Att) within ViT to guide the network to focus on the most artifact-prone patches, termed vulnerable patches. In this section, we start by recalling the concept of blending-based data synthesis that is an essential component of L2-Att. Note that FakeFormer is trained using only real data and blending-based pseudo-fakes. Then, we define vulnerable patches, and detail their extraction process as well as the generation of the ground-truth data. Finally, we describe the architecture of the proposed L2-Att module.
4.3 Background on Blending-based Data Synthesis
This pseudo-fake generation technique has been widely used in deepfake detection [37, 47]. Given two real facial images $\mathbf{I}_{B}$ and $\mathbf{I}_{F}$, it consists in generating a manipulated image denoted as $\mathbf{I}_{M}$ as follows,
$\mathbf{I}_{M}=\mathbf{M}\odot\mathbf{I}_{F}+(\mathbf{1}-\mathbf{M})\odot\mathbf{I}_{B}\text{,}$ (3)
where $\odot$ denotes the element-wise multiplication, $\mathbf{1}$ represents an all-one matrix, and $\mathbf{M}$ refers to the deformed convex hull.
4.3.1 Vulnerable Patches
$f_{2}$ FF++ CDF1 CDF2 DFD DFW DFDC
mean 97.45 96.29 94.16 95.03 81.02 78.28
max 97.67($\uparrow$0.22) 97.25($\uparrow$0.96) 94.45($\uparrow$0.29) 96.12($\uparrow$1.09) 81.74($\uparrow$0.72) 78.91($\uparrow$0.63)
Table 1: Selection of $f_{2}$. AUC (%) comparisons between different functions $f_{2}$ (Eq. (5)).
Inspired by [45], we extend the definition of vulnerable points to vulnerable patches.
Definition 1
- Vulnerable patches are the patches in an image that are the most likely to contain blending artifacts.
This definition allows tailoring the idea of local vulnerable regions to the self-attention mechanism in ViTs, which occurs at the patch level. As in [45], we propose focusing on blending artifacts that are generic across different deepfake generation methods, thereby helping to achieve good generalization capabilities to unseen deepfakes. Inspired by [45], as vulnerable patches represent the subregions that are the most impacted by the blending, we assume that they correspond to the blending regions with the highest mixture magnitude from the original images $\mathbf{I}_{B}$ and $\mathbf{I}_{F}$. To generate vulnerable patches, we leverage blending-based synthesis techniques. We believe that data synthesis algorithms are convenient to ensure generalization while significantly reducing the need for large-scale training datasets, typically needed for transformers. Hence, similar to [37], we calculate the blending boundary $\mathbf{B}=~{}({b}_{lm})_{l\in[\![1,H]\!],m\in[\![1,W]\!]}$ resulting from the combination of $\mathbf{I}_{B}$ and $\mathbf{I}_{F}$ as follows,
$\mathbf{B}=4\cdot\mathbf{M}\odot(\mathbf{1}-\mathbf{M})\text{,}$ (4)
where $\odot$ denotes the element-wise multiplication and $\mathbf{1}$ represents an all-one matrix. A higher value of $b_{rc}$ at a specific position $(r,c)\in[1,H]\times[1,W]$ indicates a more significant mixture between $\mathbf{I}_{B}$ and $\mathbf{I}_{F}$. We then apply to $\mathbf{B}$, a patching function $f_{1}$ that extracts $N$ non-overlapping patches denoted as $\tilde{\mathbf{B}}=~{}(\tilde{\mathbf{B}}_{ij})_{i,j\in[\![1,\sqrt{N}]\!]}$ of dimension $P\times P$ such that $\tilde{\mathbf{B}}=f_{1}({\mathbf{B}})$. Finally, to quantify the vulnerability of each patch, a second function $f_{2}$ operating at the patch level is defined. In summary, the vulnerable patches $\mathcal{P}$ are obtained as follows,
$\mathcal{P}=\operatorname*{arg\,max}_{(i,j)}(f_{2}(\tilde{\mathbf{B}}_{ij})),$ (5)
For the sake of simplicity, we choose $f_{2}$ as the max operation, which implies that $f=f_{2}\circ f_{1}$ corresponds to maxpooling. We also explore the mean operation in Section 5.3. In future works, we intend to investigate additional patch-based vulnerability functions. Note that $\mathcal{P}$ can also include more than one patch, since $\tilde{\mathbf{B}}$ can be maximal at several locations. This process is illustrated in Figure 4c.
4.3.2 Ground Truth Generation for L2-Att
Finally, to obtain the ground truth to be compared to the output of L2-Att denoted as $\mathbf{S}$, we generate a weighted map $\mathbf{S}_{p}$ for each element $\mathbf{p}=(p_{x},p_{y})\in\mathcal{P}$. To take into account the neighborhood patches, we use an unnormalized Gaussian distribution to calculate $\mathbf{S}_{p}$ as follows,
$\mathbf{S}_{p}(x,y)=\exp\left(-\frac{(x-p_{x})^{2}+(y-p_{y})^{2}}{2\sigma^{2}}\right),$ (6)
where $(x,y)\in[[1,\sqrt{N}]]$ represents the spatial position, and the standard deviation $\sigma$ is fixed to $1$ by default. We obtain $\mathbf{S}$ by overlaying $\{\mathbf{S}_{p}\}_{p\in\mathcal{P}}$. The ground truth generation process is also illustrated in Figure 4c. It can be noted that, for real data, $\mathbf{S}$ is set to a zero matrix.
4.3.3 Learning-based Local Attention Module (L2-Att)
By predicting the locations of vulnerable patches, L2-Att diffuses artifact-prone local information through the ViT. In particular, L2-Att first takes the patch embeddings $\mathbf{z}^{L^{\prime}}_{\neg 0}\in\mathbb{R}^{N\times D}$ as input and processes them to produce spatial features as follows,
[…]
$\mathbf{F}_{out}$ $=\text{Reshape}(\mathbf{F}),\hskip 2.84526pt\mathbf{F}_{out}\in\mathbb{R}^{D\times\sqrt{N}\times\sqrt{N}}\text{,}$
After that, $\mathbf{F}_{out}$ is fed into two convolution blocks (ConvBlock). The predicted weighted heatmap denoted as $\hat{\mathbf{S}}$ describing the presence probability of vulnerable patches is obtained as follows,
$\hat{\mathbf{S}}=\text{ConvBlock}_{1\times 1}(\text{ConvBlock}_{3\times 3}(\mathbf{F}_{out})),$ (8)
where $\hat{\mathbf{S}}\in\mathbb{R}^{1\times\sqrt{N}\times\sqrt{N}}$. Each convolution block comprises a convolution layer, followed by a BatchNorm and a RELU layer. The notation $\text{ConvBlock}_{n\times n}$ indicates a convolution block with an ${n\times n}$ filter size. A detailed illustration of the L2-Att module can be seen in Figure 4b.
Method Training data Params FLOPs Test set AUC(%)
Real Fake FF++ CDF2
ViT-B [60] $\checkmark$ $\checkmark$ 84M 55.4G - 82.33
ViT-B+TALL [60] $\checkmark$ $\checkmark$ 84M 55.4G - 86.58
FakeFormer-S $\checkmark$ 23M 8.9G 97.67 94.45
FakeFormer-B $\checkmark$ 91M 35.8G 97.76 95.21
Swin-B [60] $\checkmark$ $\checkmark$ 86M 47.5G - 83.13
Swin-B+TALL [60] $\checkmark$ $\checkmark$ 86M 47.5G 99.87 90.79
FakeSwin-S $\checkmark$ 55M 6.5G 99.89 94.48
FakeSwin-B $\checkmark$ 91M 11.5G 99.94 95.44
Table 2: Performance of Transformer-based approaches. Comparisons in terms of AUC (%) on CDF2 [38], and FF++ [46]. Methods are trained on FF++ [46] and tested on the other datasets. Bold and underlined results highlight the best and the second-best performance, respectively.
Method Fake Test set (%)
CDF1 CDF2 DFW DFD DFDCP DFDC
AUC AP AUC AP AUC AP AUC AP AUC AP AUC AP
Xception [46] $\checkmark$ 58.81 65.59 61.18 66.93 65.29 55.37 89.75 85.48 69.90 91.98 45.60 -
FaceXRay [37] $\checkmark$ 80.58 73.33 - - - - 95.40 93.34 80.92 72.65 - -
LocalRL [13] $\checkmark$ - - 78.26 - - - 89.24 - 76.53 - - -
Multi-attentional [64] $\checkmark$ 69.14 74.03 68.26 75.25 73.56 73.79 92.95 96.51 83.81 96.52 - -
RECCE [7] $\checkmark$ 49.96 63.04 70.93 70.35 68.16 54.41 98.26 79.42 80.93 92.76 - -
SFDG [57] $\checkmark$ - - 75.83 - 69.27 - 88.00 - 73.63 - - -
CADDM [21] $\checkmark$ 89.36 93.25 93.88 91.12 74.48 75.23 99.03 99.59 - - 73.85 -
UCF [62] $\checkmark$ - - 82.4 - - - 94.5 - 80.5 - - -
M2TR [55] $\checkmark$ - - 68.2 - - - - - - - - -
DFDT [33] $\checkmark$ - - 88.3 - - - - - 76.1 - - -
TALL-Swin [60] $\checkmark$ - - 90.79 - - - - - 76.78 - - -
ICT [22] 81.43 - 85.71 - - - 84.13 - - - - -
SBI [47] 92.53 79.91 93.18 85.16 67.47 55.87 97.56 92.79 86.15 93.24 72.42 -
FakeFormer 97.25 98.36 94.45 97.15 81.74 83.72 96.12 98.31 96.30 99.50 78.91 80.01
Table 3: Cross-dataset evaluation. Comparisons in terms of AUC (%) and AP (%) on CDF1 [38], CDF2 [38], DFW [65], DFD [24], DFDCP [20], and DFDC [19]. ✓symbol is used to depict methods that utilized both Real data and Fake data for training.
4.4 Training Objective
To train FakeFormer, we optimize two losses, namely the Binary Cross Entropy (BCE) loss for classification denoted as $L_{cls}(\hat{\mathbf{y}},\mathbf{y})$, and the regression loss related to the prediction of vulnerable patches locations, denoted as $L_{att}(\hat{\mathbf{S}},\mathbf{S})$. Therefore, the total loss $L$ is defined as follows,
$L=L_{cls}+\lambda L_{att}\text{,}$ (9)
where $\lambda$ is a balancing factor between the two losses and is empirically set to $10$. We employ the focal loss [39] to compute $L_{att}(\hat{\mathbf{S}},\mathbf{S})$.
5 Experiments
5.1 Setups
Datasets. As in [37, 47, 26, 12, 11, 21, 60], we use the FaceForensics++ (FF++) [46] dataset for training. Only real data and blended-based pseudo-fakes [37, 47] are employed for training. We consider both the in-dataset and the cross-dataset protocols by testing on FF++ and other benchmarks including Celeb-DF (CDF1, CDF2) [38], WildDeepfake (DFW) [65], DeepFakeDetection (DFD) [24], Deepfake Detection Challenge Preview (DFDCP) [20], and Deepfake Detection Challenge (DFDC) [19], respectively. More details about the datasets are provided in the supplementary materials.
Data Processing. Following the splitting convention [46], we sample $128$, $32$, and $32$ frames from each video for training, validation, and testing respectively. This significantly reduces the amount of training data compared to previous ViT-based works [22, 33, 55]. The facial regions are extracted using RetinaFace [18]. To align two images for generating pseudo-fakes [37, 47], Dlib [34] is used as a landmark extractor. Further details are provided in the supplementary materials.
Training. We train FakeFormer for $200$ epochs using the AdamW [41] optimizer with a weight decay of $10^{-4}$ and a batch size of $32$. The weights are initialized using a ViT [9, 23] pretrained on ImageNet [17]. The learning rate maintains at $5\times 10^{-5}$ during the first quarter of iterations, then gradually decays to zero over the remaining epochs. We freeze the backbone (i.e., ViT without the head) for the first $6$ epochs, before training all layers. As a regularizer, we integrate LabelSmoothing [44] with the loss function (Eq. 9). For each video in the batch data, we dynamically select only $m$ frames, with $m=8$ or $m=16$ when using SBI [47] or BI [37], respectively. Data augmentation techniques include color jittering, random cropping, scaling, horizontal flipping, Gaussian noise, blurring, and JPEG compression. All experiments are conducted on $4$ NVIDIA A100 GPUs.
Method Params Test set AUC(%)
CDF2 DFDCP DFDC
FaceXRay+BI [37] 41.23M 79.5 65.5 -
ICT [22]+BI [37] 21.45M 85.71 - -
LAA-Net [45]+BI [37] 27.13M 86.28 69.69 64.18
FakeFormer+BI [37] 22.77M 90.34 78.71 76.84
Table 4: Performance using BI [37]. Comparisons in terms of AUC (%) on CDF2 [38], DFDCP [20], and DFDC [19].
Evaluation Metrics. To compare FakeFormer with existing works, we employ two common evaluation metrics: the Area under the Curve (AUC) and the Average Precision (AP). For a fair comparison, similar to previous studies [21, 63, 47, 4, 60, 28], we report the video-level performance.
Method Params FLOPs Test set AUC(%)
CDF2 DFDC DFDCP
SBI [47] 19.34M 8.3G 93.18 72.42 86.15
LAA-Net [45] 27.13M 11.6G 95.40 72.43 86.94
FakeFormer 22.77M 8.9G 94.45 78.91 96.30
Table 5: Performance using SBI [47]. Comparisons in terms of AUC (%) on CDF2 [38], DFDCP [20], and DFDC [19].
Architectures. We utilize two variants of FakeFormer that we call FakeFormer-S and FakeFormer-B. By default, we use the lightweight FakeFormer-S, where $H=W=112$ and $P=8$. FakeFormer is based on the vanilla vision transformer, i.e., ViT [23]. However, although FakeFormer is based on ViT, we also assess its applicability using another transformer architecture, namely FakeSwin, which is based on Swin [40]. Similarly to FakeFormer, we consider two variants: FakeSwin-S and FakeSwin-B. Further details are provided in the supplementary materials.
Method Fake Saturation Block Noise Blur Avg
FaceXray [37] $\checkmark$ 97.6 99.1 49.8 63.8 77.6
LipForensics [28] $\checkmark$ 99.9 87.4 73.8 96.1 89.3
LAA-Net [45] 99.9 99.9 53.9 98.2 87.9
FakeFormer 98.1 97.0 75.3 87.4 89.5
FakeSwin 99.8 99.9 81.6 89.8 92.8
Table 6: Robustness to perturbations on FF++ . Focusing on vulnerable patches, FakeFormer shows improved robustness to common perturbations, especially to noise.
5.2 Comparison with State-of-the-art
In this section, we compare the performance of FakeFormer to SOTA methods. In all tables, bold and underlined results highlight the best and the second-best performance, respectively.
Model L2-Att FF++ CDF1 CDF2 DFD DFDC Avg.
ViT [23] 97.48 94.42 92.62 95.72 77.35 91.52
FakeFormer $\checkmark$ 97.67($\uparrow$0.19) 97.25($\uparrow$2.83) 94.45($\uparrow$1.83) 96.12($\uparrow$0.40) 78.91($\uparrow$1.56) 92.88($\uparrow$1.36)
Swin [40] 99.75 94.46 90.89 99.59 74.20 91.78
FakeSwin $\checkmark$ 99.89($\uparrow$0.14) 96.69($\uparrow$2.23) 94.48($\uparrow$3.59) 99.68($\uparrow$0.09) 77.47($\uparrow$3.27) 93.64($\uparrow$1.86)
Table 7: L2-Att ablation. Performance of FakeFormer without and with L2-Att using the cross-dataset setting.
Comparison with Transformer-based Approaches. Table 2 compares FakeFormer to Transformer-based methods in terms of model size (Params), computational cost (FLOPs), and AUC on FF++ and CDF2. Even with the smaller variant (i.e., -S), our method outperforms vanilla ViT-B and Swin-B. This highlights the effectiveness of the proposed L2-Att module, resulting in a lightweight model. Moreover, regardless of the transformer backbone, FakeFormer and FakeSwin generalize better than TALL [60] despite their simplicity. It should be noted that the ViT-B and the Swin-B results are directly recovered from [60], which means that the backbone configurations might be different.
Model Target Params FLOPs Test set AUC(%)
CDF1 CDF2 DFD DFDC Avg.
FakeFormer VPoint 23.61M 9.4G 97.22 93.39 93.70 77.71 90.51
VPatch 22.77M 8.9G 97.25 94.45 96.12 78.91 91.68($\uparrow$1.17)
FakeSwin VPoint 57.25M 8.1G 96.16 94.25 99.30 74.99 91.18
VPatch 54.89M 6.5G 96.69 94.48 99.68 77.47 92.08($\uparrow$0.9)
Table 8: VPoint versus VPatch. AUC (%) when using Vulnerable Points [45] (VPoint) versus Vulnerable Patches (VPatch).
Blending-based Synthesis Approaches. Table 4 and Table 5 compare our method to SOTA when using Blending-Images (BI) [37] and Self-Blended-Images (SBI) [47], respectively. As observed, we generally achieve better performances than SOTA methods using the same blending algorithm on CDF2, DFDC, and DFDCP datasets, despite training on a drastically lower number of data as compared to other ViT-based approaches such as ICT [22]. Moreover, LAA-Net [45], as a CNN-based method, requires a larger model size as compared to our lightweight FakeFormer.
Overall Cross-dataset Evaluation. The overall comparison with SOTA methods under the cross-dataset setup can be seen in Table 3. As observed, FakeFormer generalizes better than SOTA methods, including those trained with fake data, those trained with pseudo-fakes (without fake data), and transformer-based models such as ICT [22], M2TR [55], and DFDT [33], although our model is trained with a significantly lower amount of data.
$\lambda$ CDF1 CDF2 DFD DFDC Avg.
1 94.68 93.67 94.62 77.91 90.22
10 97.25 94.45 96.12 78.91 91.68
100 97.86 94.96 94.88 78.58 91.57
Table 9: Variation of the parameter $\lambda$. AUC (%) comparison using different $\lambda$ values (Eq. (9)).
Robustness to Perturbations. Since deepfakes can be altered by common perturbations on social media. Following the settings in [45, 28], experiments evaluating the robustness of FakeFormer against perturbations on FF++ [46] are reported in Table 6. We can observe the superiority of FakeFormer and FakeSwin as compared to existing methods.
5.3 Ablation Study
Impact of L2-Att. To validate the L2-Att module, we compare ViT [23]/Swin [40] and FakeFormer (ViT+L2-Att)/FakeSwin (Swin+L2-Att) (with SBI [47]). The results on several datasets [46, 38, 24, 19] are presented in Table 7. In general, L2-Att consistently contributes to the enhancement of both ViT and Swin, confirming the relevance of the proposed explicit attention mechanism.
Vulnerable Points versus Vulnerable Patches. LAA-Net [45] proposes an attention mechanism based on vulnerable points (VPoint). To demonstrate the compatibility of vulnerable patches (VPatch) with transformers, we report in Table 8 the obtained results when replacing VPatch with VPoint within FakeFormer and FakeSwin. We note that the use of VPatch not only results in better performance but also guarantees a lower computational cost as compared to VPoint. We note that the higher number of parameters and FLOPs associated with using VPoint is caused by the decoder designed to locate VPoint. Meanwhile, VPatch does not require any decoder, making it more computationally efficient.
Selection of $f_{2}$. Table 1 compares two aggregation functions $f_{2}$ defined in Eq. (5) coupled with FakeFormer: mean and max operations. In both cases, the stability of the results can be seen. By default, we select the max operation as it gives slightly better results.
Parameter $\lambda$. The weight $\lambda$ defined in Eq. (9) is set empirically to $10$ as it yields the best performance on average. We report the results using different values of $\lambda$ within FakeFormer in Table 9. It can be observed the generalization across four testing benchmarks remains robust regardless of the value of $\lambda$.
6 Limitations
As FakeFormer is tailored to blending-based manipulations such as face-swaps and face reenactments, it shows poor performance when tested on fully synthetic data. Indeed, we obtain $50.84$% and $59.17$% AUC when testing FakeFormer and FakeSwin on $1000$ randomly generated images using stylegan2 [32].
7 Conclusion
In this paper, we identify the shortcomings of ViTs in capturing localized features, explaining its restricted adoption in the field of deepfake detection. To address this, we introduce FakeFormer, a lightweight ViT-based framework for generalizable deepfake detection. FakeFormer incorporates L2-Att, an attention strategy directing the focus toward local inconsistency-prone regions, thus enhancing the detection capabilities of transformers. Leveraging the synthetic data generation and the proposed attention mechanism mitigates the need for extensive training data and large models. Detailed experiments demonstrate the superiority of FakeFormer over state-of-the-art methods. In future work, we will investigate strategies to extend FakeFormer for capturing temporal information and imposing more robustness to fully synthetic deepfakes.
8 Appendix
8.1 Code Release
The code is provided as a part of the supplementary materials. The code is under license registration and will be publicly released soon with pretrained weights.
8.2 More Details regarding the Datasets
Datasets. FF++ contains $1000$ original videos and $4000$ deepfake videos. The types of deepfakes include Deepfakes (DF) [48], Face2Face (F2F) [52], FaceSwap (FS) [35], and NeuralTextures (NT) [51]. Moreover, while CDF and DFD represent datasets including high-quality deepfakes, DFW stands out as one of the most challenging benchmarks. DFW is collected entirely from the internet without any prior knowledge regarding the used manipulation techniques. Additionally, it includes images with occlusions, illumination changes, different head poses, and diverse backgrounds. Meanwhile, DFDC and its preview version called DFDCP, comprise more than $128,000$ videos and is therefore one of the largest datasets in the field of deepfake detection.
Data Processing. The RetinaFace [18] algorithm is employed to extract facial bounding boxes. These boxes are enlarged by a factor of $1.25$ and then resized to a specific resolution $(H,W)$ to match the different backbone configurations. To generate pseudo-fakes, the two images are first aligned. For this purpose, the Dlib [34] library is utilized to extract $68$ and $81$ facial landmark for the two considered blending synthesis algorithms, BI [37] and SBI [47], respectively. Finally, the processed pseudo-fakes are generated for each training iteration.
Architecture Details. Here, we describe the hyperparameters of two considered FakeFormer variants:
  • FakeFormer-S: $H=W=112$, $P=8$, $L=12$, $D=384$, MLP size $=1536$, No. Heads $=6$, Params$=23$M, FLOPs$=8.9$G.
  • FakeFormer-B: $H=W=224$, $P=16$, $L=12$, $D=768$, MLP size $=3072$, No. Heads $=12$, Params$=91$M, FLOPs$=35.8$G.
where the MLP size represents the dimension of hidden layers in MLP, the No. Heads denotes the number of heads in MHSA, Params is the number of parameters, and FLOPs represents the computational cost in terms of floating point operations per second. And for FakeSwin architecture, we adopt these two backbone variants from Swin [40], namely:
  • FakeSwin-S: $H=W=224$, $P=4$, $M=7$, $d=32$, $\alpha=4$, $C_{h}=96$, Layer Numbers = {2, 2, 18, 2}, No. Heads = {3, 6, 12, 24}, Params$=55$M, FLOPs$=6.5$G.
  • FakeSwin-B: $H=W=224$, $P=4$, $M=7$, $d=32$, $\alpha=4$, $C_{h}=128$, Layer Numbers = {2, 2, 18, 2}, No. Heads = {4, 8, 16, 32}, Params$=91$M, FLOPs$=11.5$G.
where $M$ is the window size, $d$ is the query dimension of each head, the expansion layer of each MLP is $\alpha$, and $C_{h}$ denotes the number channel of the hidden layers in the first stage.
8.3 More Details of training Setups in Section 3
In this section, we present more details of the experiment settings in Section 3.
Specifically, in Figure 3a, we train vanilla ViT [23] and Swin [40] with SBI data synthesis [47] using the same training settings described in [45, 47]. Besides, the detection performance of three CNN methods in different ranges of Mask-SSIM [42] on CDF2 [38], i.e., LAA-Net [45], CADDM [21], and SBI [47] for the comparison is directly extracted from [45].
Secondly, in Figure 3b, all models including CNNs and variants of ViTs are trained on FF++ [46] with both real and fake data for $50$ epochs. Following conventional spitting [46], we uniformly extract $128$ and $32$ frames of each video for training and validation, respectively. Hence, there are in total of $460800$ and $22400$ frames for the corresponding task. The weights of models are initialized by pretrained on ImageNet [17]. We employ Adam for CNNs and AdamW for ViTs as optimizers. The learning rate is initially set to $10^{-4}$ and linearly decays to 00 at the end of the training period. All experiments are carried out using a NVIDIA A100 GPU.